OpenStudy (anonymous):

If the eccentricity of a planet's orbit about the Sun is 0.4, find the ratio of the lengths of the major to the minor axes of the orbit of the planet.

OpenStudy (yrelhan4):

planet's orbit around the sun is an elipse.. whats the formula for eccentricity of an elipse?

OpenStudy (anonymous):

\[\epsilon =\sqrt{1-\frac{ b ^{2} }{ a ^{2} }}\]

OpenStudy (yrelhan4):

true. now you just have to solve it.. put 0.4 for e, and solve for a/b..

OpenStudy (anonymous):

now I got it \[\frac{ b ^{2} }{ a ^{2} }=1-\epsilon ^{2}=1-0.16=0.84\] \[\frac{ a }{ b }=\sqrt{\frac{ 1 }{ 0.84 }}\] \[\frac{ a }{ b }=1.091\] Thanks.

OpenStudy (yrelhan4):

Cheers. :)